# X-Vine Models for Multivariate Extremes

### v1

@@ 5 Vine decompositions of tail copula densities
$r(\bm{x})=\prod_{e\in E_{1}}r_{a_{e},b_{e}}(x_{a_{e}},x_{b_{e}})\cdot\prod_{j=2}^{q}\prod_{e\in E_{j}}c_{a_{e},b_{e};D_{e}}\left({R_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}}),R_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}})}\right),$
with an empty second product in case $q=1$. A truncated X-vine specification involves $d-1$ bivariate tail copula densities and $\sum_{j=2}^{q}(d-j)$ bivariate copula densities, yielding $qd-q(q+1)/2$ bivariate components in total. The truncation level $q$ allows to tune the trade-off between sparsity and flexibility. If $q=1$, the model is a Markov tree as in Engelke and Hitz (2020) and Segers (2020). Increasing $q$ and adding trees (‘layers’) yield more complex dependence structures. Truncated X-vine specifications will be shown at work in the case study in Section 9.
@@ 6.1 Inverted multivariate Pareto distributions
The $d$-variate random vector $\bm{Z}$ has an inverted multivariate Pareto distribution if there exists a $d$-variate tail copula measure $R$ such that
$\operatorname{\mathsf{P}}(\bm{Z}\in B)=\frac{R(B\cap\mathbb{L})}{R(\mathbb{L})},$
for Borel sets $B$, with $\mathbb{L}=\left\{{\bm{x}\in(0,\infty]^{d}:\min\bm{x}<1}\right\}$; equivalently, if there exists a multivariate Pareto random vector $\bm{Y}$ such that $\bm{Z}=1/\bm{Y}$.
If $R$ is concentrated on $(0,\infty)^{d}$ and has density $r$, then $\bm{Z}$ has probability density $\bm{x}\mapsto r(\bm{x})\operatorname{\mathds{1}}(\bm{x}\in\mathbb{L})/R(\mathbb{L})$. In general, for $j\in[d]$, the conditional distribution of $Z_{j}$ given $Z_{j}<1$ is uniform on $(0,1)$; this is a consequence of the marginal constraint (2.2). More generally, for Borel sets $B$, we have
@@
For non-empty $J\subset[d]$, we have
$\left({t\bm{U}_{J}\mid\min\bm{U}_{J}<1/t}\right)\rightsquigarrow\bm{Z}_{|J}\stackrel{{\scriptstyle d}}{{=}}\left({\bm{Z}_{J}\mid\min\bm{Z}_{J}<1}\right),\qquad t\to\infty.$
The conditional marginal $\bm{Z}_{|J}$ has a $|J|$-variate inverted multivariate Pareto distribution associated with the marginal tail copula measure $R_{J}$ in (2.6). In contrast, the random vector $\bm{Z}_{J}$ does not necessarily have an inverted multivariate Pareto distribution, as $\min\bm{Z}_{J}<1$ is not guaranteed.
6.2 Random sampling from (inverted) multivariate Pareto distributions
Through equations (2.10) and (6.2), the (inverted) multivariate Pareto distribution serves as a model for a random vector conditionally on the event that at least one variable takes a value far in the tail of its respective marginal distribution. The L-shaped support of the (inverted) multivariate Pareto distribution makes direct random sampling from it a little awkward. Lemma 2 in Engelke and Hitz (2020) provides an ingenious algorithm that reduces the task of simulating from a multivariate Pareto random vector $\bm{Y}$ to random sampling from the conditional distributions $\left({\bm{Y}\mid Y_{j}>1}\right)$ for every $j\in[d]$. Below, we will study the equivalent problem of simulating from the conditional distribution of $\left({\bm{Z}\mid Z_{j}<1}\right)$ in case the tail copula density is an X-vine as in Definition 5.4.
We will do so by inverting the Rosenblatt transformation (Rosenblatt, 1952), applying conditional quantile functions with an increasing number of conditioning variables successively to independent uniform random variables. A judicious choice of the ordering of the variables permits to compute the required conditional quantile functions recursively in terms of the bivariate ingredients of the X-vine specification. This ordering is encoded in the permutation constructed in the next lemma.
@@ 6.2 Random sampling from (inverted) multivariate Pareto distributions
Let $\mathcal{V}$ be a regular vine sequence on $d$ elements. For every $j\in[d]$, there exists a permutation $\sigma_{j}$ of $\left\{{1,\ldots,d}\right\}$ such that (i) $\sigma_{j}(1)=j$ and (ii) there exist edges $e_{j,1}\in E_{1},\ldots,e_{j,d-1}\in E_{d-1}$ such that $\sigma_{j}(k)\in\mathcal{C}_{e_{k-1}}$ and $\left\{{\sigma_{j}(i):i=1,\ldots,k}\right\}=A_{e_{j,k-1}}$ for each $k\in\left\{{2,\ldots,d}\right\}$.
Let $r$ be an X-vine tail copula density. For each edge $e\in E_{2}\cup\ldots\cup E_{d-1}$ and for fixed $10<u_{b_{e}}<$ 1, let $u_{a_{e}}\mapsto C_{a_{e}|b_{e};D_{e}}^{-1}\left({u_{a_{e}}|u_{b_{e}}}\right)$ be the inverse of the distribution function $u_{a_{e}}\mapsto C_{a_{e}|b_{e};D_{e}}\left({u_{a_{e}}|u_{b_{e}}}\right)$, with $C_{a_{e}|b_{e};D_{e}}$ defined in (5.3). Similarly for $C_{b_{e}|a_{e};D_{e}}^{-1}$. Recall from Proposition 3.1 that for $i\in[d]$, non-empty $J\subseteq[d]\setminus\left\{{i}\right\}$, and $\bm{x}_{J}\in(0,\infty)^{d}$ such that $r_{J}(\bm{x}_{J})>0$, the quantile function $u_{i}\mapsto R^{-1}_{i|J}(u_{i}|\bm{x}_{J})$ is the inverse of the distribution function $x_{i}\mapsto R_{i|J}(x_{i}|\bm{x}_{J})$. Inverting (5.4) yields the recursive relations
@@
$\begin{split}Z_{j}^{(j)}&=W_{j},\\ Z_{\sigma_{j}(k)}^{(j)}&=R^{-1}_{\sigma_{j}(k)|\sigma_{j}(\left\{{1,\ldots,k-1}\right\})}\left({W_{\sigma_{j}(k)}\mid\bm{Z}_{\sigma_{j}(\left\{{1,\ldots,k-1}\right\})}^{(j)}}\right),\qquad k\in\left\{{2,\ldots,d}\right\}.\end{split}$ (6.4)
@@
The actual simulation algorithm is based on structure matrices encoding regular vine sequences as explained in Appendix A. In Section 8, we will apply the simulation technique based on Proposition 6.3 to assess the estimation methods we will introduce in Section 7 through Monte Carlo experiments.
@@ 7 Estimation and model selection
Let $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,d})$ for $i\in\left\{{1,\ldots,n}\right\}$ be an independent random sample from a distribution function $F$ with continuous but unspecified margins $F_{1},\ldots,F_{d}$ and whose survival copula $C$ has lower tail copula $R$ (Definition 2.1). Suppose that the tail copula density $r$ is an X-vine with specification $\left({\mathcal{V},\mathcal{R},\mathcal{C}}\right)$ (Definitions 5.4 and 5.6 and Theorem 5.7). We propose a procedure to estimate $r$ from the excesses over a high multivariate threshold.
@@ 7.1 Copulas and inverted multivariate Pareto distributions
Let $r$ be a $d$-variate copula tail copula density, not necessarily an X-vine. In Sklar’s theorem (Proposition 3.1), suppose that the copula density $c_{I;J}$ satisfies the simplifying assumption (Definition 3.4). The following proposition shows how to transform an inverted multivariate Pareto random vector $\bm{Z}$ associated to $r$ into a random vector with density $c_{I;J}$.
@@
By statement (ii), the density of $\left({R_{i|J}(Z_{i}|\bm{Z}_{J})}\right)_{i\in I}$ given $\bm{Z}_{J}\in A$ is equal to $c_{I;J}$ for any non-empty set $A\subseteq\left\{{\bm{z}\in(0,\infty)^{J}:\min\bm{z}<1}\right\}$. For estimation, we will use this property for $A=(0,1)^{J}$, requiring in effect that all variables (rather than at least one) in $J$ exceed a high threshold. We do so in order to avoid a potential bias stemming from including too many non-extreme values in the procedure. An alternative would be to opt for a censored likelihood approach (Ledford and Tawn, 1996), but this seems difficult to implement beyond dimensions two or three.
@@ 7.2 Sequential maximum likelihood estimation of X-vines
Let the $d$-variate tail copula density $r$ be an X-vine specified by $(\mathcal{V},\mathcal{R},\mathcal{C})$ as in Theorem 5.7. Assume that the bivariate tail copula densities ($\mathcal{R}$) and bivariate copula densities ($\mathcal{C}$) belong to specified parametric families. Let $\bm{\theta}=(\theta_{\mathcal{R}},\theta_{\mathcal{C}})$ denote the parameter vector associated with $r$. Here, $\theta_{\mathcal{R}}=(\theta_{e})_{e\in E_{1}}$ contains the parameters (or parameter vectors) $\theta_{e}\in\Theta_{e}$ associated with each pairwise tail copula density $r_{a_{e},b_{e}}$ for $e\in E_{1}$, while $\theta_{\mathcal{C}}=\left({\theta_{\mathcal{C},j}}\right)_{j=2}^{d-1}$ for $\theta_{\mathcal{C},j}=(\theta_{e})_{e\in E_{j}}$ denotes the vector of parameters (or parameter vectors) $\theta_{e}\in\Theta_{e}$ associated with each bivariate copula density $c_{a_{e},b_{e};D_{e}}$ for $e\in\bigcup_{j=2}^{d-1}E_{j}$. While it is possible to derive the full-likelihood of an X-vine (inverted) multivariate Pareto distribution, performing parameter estimation with the full model in high dimensions becomes challenging in extremes. Instead, using the X-vine decomposition into bivariate components and recursively defined quantities (Theorem 5.7), we outline a sequential procedure for parameter estimation, starting from $T_{1}$ to $T_{d-1}$ tree by tree. This approach is inspired by the one for regular vine copulas (see, e.g., Czado, 2019), but with suitable adaptations to the extreme value context.
(1) Standardising the margins and selecting sub-samples
Recall from the beginning of the section that $\bm{X}_{1},\ldots,\bm{X}_{n}$, with $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,d})$, is an independent random sample from the unknown $d$-variate distribution $F$, whose survival copula $C$ is assumed to have tail copula $R$ with tail copula density $r$. Set
$\widehat{U}_{i,j}=1-\widehat{F}_{j}(X_{i,j}),\qquad i=1,\ldots,n,\ j\in[d],$ (7.1)
where $\widehat{F}_{j}$ denotes any estimator of the marginal distribution function $F_{j}(x)=\operatorname{\mathsf{P}}(X_{i,j}\leqslant x)$. One possibility is the empirical distribution function, and to avoid boundary effects, we set
@@
By Eq. (6.2), for large $t>0$, the rescaled points $t\widehat{\bm{U}}_{i}$ for $i\in\left\{{1,\ldots,n}\right\}$ such that $\min\widehat{\bm{U}}_{i}<1/t$ constitute pseudo-observations from a distribution that approximates the inverted multivariate Pareto distribution associated with $r$. We set $t=n/k$ where $k\in\left\{{1,\ldots,n}\right\}$ is such that $k$ is ‘large’ and $k/n$ is ‘small’; in the usual large-sample setting in extreme value theory, we have $k=k_{n}$ with $k\to\infty$ and $k/n\to 0$. For $j\in[d]$, let
@@
be the set of indices $i$ corresponding to observations whose the $j$th component $X_{i,j}$ is ‘large’; for the rank-based definition of $\widehat{U}_{i,j}$ in the previous paragraph, this set of indices corresponds to the $k$ largest observations for the $j$th variable. Further, put
$K_{J}=\bigcap_{j\in J}K_{j}\qquad\text{and}\qquad K=\bigcup_{j\in[d]}K_{j},$
which are, respectively the set of indices $i$ corresponding to large observations in all variables in $J\subseteq[d]$ simultaneously and the set of indices with large observations in at least one variable. Write $\widehat{\bm{Z}}_{i}=(\widehat{Z}_{i,j})_{j\in[d]}$ where $\widehat{Z}_{i,j}=\left({n/k}\right)\widehat{U}_{i,j}$. In view of Eq. (6.2), we treat $\{\widehat{\bm{Z}}_{i}\}_{i\in K}$ as a sample of $|K|$ pseudo-observations of the inverted multivariate Pareto distribution associated with $r$. The sample size $|K|$ is random, and in the rank-based definition above, we have $k\leqslant|K|\leqslant dk$. For non-empty $J\subseteq[d]$ and $i\in K$, we write $\widehat{\bm{Z}}_{i,J}=(\widehat{Z}_{i,j})_{j\in J}$.
(2) Estimating the tail copula parameters $\theta_{\mathcal{R}}$ in $T_{1}$
For each edge $e=\left\{{a_{e},b_{e}}\right\}\in E_{1}$, we estimate the parameter (vector) $\theta_{e}$ associated with the bivariate tail copula density $r_{a_{e},b_{e}}(\,\cdot\,;\theta_{e})$. By Eq. (6.1), for $j\in\left\{{a_{e},b_{e}}\right\}$, the conditional density of $(Z_{a_{e}},Z_{b_{e}})$ given $Z_{j}<1$ is $r_{a_{e},b_{e}}(z_{a_{e}},z_{b_{e}};\theta_{e})\operatorname{\mathds{1}}(z_{j}<1)$. As one out of many possibilities, we use maximum pseudo-likelihood estimation to fit this density to the sub-samples $(\widehat{Z}_{i,a_{e}},\widehat{Z}_{i,b_{e}})$ for both $i\in K_{a_{e}}$ and $i\in K_{b_{e}}$, with $K_{j}$ as in Eq. (7.3), More precisely, we maximise each of the two pseudo-likelihoods
@@
over $\theta_{e}^{(j)}\in\Theta_{e}$, yielding estimates $\hat{\theta}_{e}^{(a_{e})}$ and $\hat{\theta}_{e}^{(b_{e})}$, respectively. The final estimate is $\hat{\theta}_{a_{e},b_{e}}=\{\hat{\theta}_{e}^{(a_{e})}+\hat{\theta}_{e}^{(b_{e})}\}/2$. The idea of averaging maximum pseudo-likelihood estimators on product spaces has already been suggested for the Hüsler–Reiss model in (Engelke et al., 2015), and implemented for multivariate Pareto distributions in the graphicalExtremes package (Engelke et al., 2022). Other approaches for parameter estimation can be considered: censored likelihoods (Ledford and Tawn, 1996; de Haan et al., 2008) or the empirical stable tail dependence functions (Einmahl et al., 2008, 2018).
(3) Estimating the copula parameters $\theta_{\mathcal{C}}$ in $T_{2},\ldots,T_{d-1}$
Start with an edge $e=(a_{e},b_{e};D_{e})\in E_{2}$, i.e., such that $D_{e}$ is a singleton, $D_{e}=\left\{{d_{e}}\right\}$, say. Then $f=\left\{{a_{e},d_{e}}\right\}$ and $g=\left\{{b_{e},d_{e}}\right\}$ are edges in $E_{1}$. By Proposition 7.1(ii), the density of the random pair
$\left({R_{a_{e}|d_{e}}(Z_{a_{e}}|Z_{d_{e}};\theta_{f}),R_{b_{e}|d_{e}}(Z_{b_{e}}|Z_{d_{e}};\theta_{g})}\right)$
is $c_{a_{e},b_{e};D_{e}}(\,\cdot\,;\theta_{e})$. The tail copula density parameters $\theta_{f}$ and $\theta_{g}$ have been estimated in the previous step, yielding estimates $\hat{\theta}_{f}$ and $\hat{\theta}_{g}$. Define pseudo-observations $(\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}})$ for $i\in K_{D_{e}}=K_{d_{e}}$ from $c_{a_{e},b_{e};D_{e}}$ by
$\widehat{U}_{i,a_{e};D_{e}}$ $=R_{a_{e}|D_{e}}\left({\widehat{Z}_{i,a_{e}}\mid\widehat{Z}_{i,d_{e}};\;\hat{\theta}_{f}}\right),$ $\widehat{U}_{i,b_{e};D_{e}}$ $=R_{b_{e}|D_{e}}\left({\widehat{Z}_{i,b_{e}}\mid\widehat{Z}_{i,d_{e}};\;\hat{\theta}_{g}}\right).$
We estimate the parameter (vector) $\theta_{e}\in\Theta_{e}$ by maximising the pseudo-likelihood
$\mathcal{L}_{\mathcal{C},e}\left({\theta_{e};\;\left({\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}}}\right),i\in K_{D_{e}}}\right)=\prod_{i\in K_{D_{e}}}c_{a_{e},b_{e};D_{e}}\left({\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}};\,\theta_{e}}\right).$ (7.5)
Other estimators are possible based on parametric relationships between the dependence coefficients such as Kendall’s tau and Spearman’s rho in relation to the parameters of corresponding bivariate copula densities (Czado, 2019).
Parameters (or parameter vectors) associated with edges $e=(a_{e},b_{e};D_{e})$ in $E_{j}$ for layers $j\in\left\{{3,\ldots,d-1}\right\}$ can be estimated similarly, proceeding recursively. Suppose that all parameters associated with edges in $E_{1:(j-1)}=E_{1}\cup\cdots\cup E_{j-1}$ have already been estimated; let $\theta(E_{1:(j-1)})$ denote the parameters and $\hat{\theta}(E_{1:(j-1)})$ their estimate. By Proposition 7.1(i), the density of the random pair
$\left({R_{a_{e}|D_{e}}\left({Z_{a_{e}}\mid\bm{Z}_{D_{e}};\;\theta(E_{1:j-1})}\right),\;R\left({Z_{b_{e}}\mid\bm{Z}_{D_{e}};\;\theta(E_{1:j-1})}\right)}\right)$
is $c_{a_{e},b_{e};D_{e}}$. Define pseudo-observations $(\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}})$ for $i\in K_{D_{e}}$ from $c_{a_{e},b_{e};D_{e}}$ by
$\widehat{U}_{i,a_{e};D_{e}}$ $=R_{a_{e}|D_{e}}\left({\widehat{Z}_{i,a_{e}}\mid\widehat{\bm{Z}}_{i,D_{e}};\;\hat{\theta}(E_{1:j-1})}\right),$ $\widehat{U}_{i,b_{e};D_{e}}$ $=R_{b_{e}|D_{e}}\left({\widehat{Z}_{i,b_{e}}\mid\widehat{\bm{Z}}_{i,D_{e}};\,\hat{\theta}(E_{1:j-1})}\right).$ (7.6)
In this definition, the conditional distribution functions $R_{a_{e}|D_{e}}$ and $R_{b_{e}|D_{e}}$ are calculated recursively via Eq. (5.6) with $e$ in that formula replaced by the edges $f,g\in E_{j-1}$ that are joined by $e$, that is, such that $e=\left\{{f,g}\right\}$. The parameter (vector) $\theta_{e}$ is then obtained by maximising the pseudo-likelihood in (7.5).

### v2

@@ 5.3 Truncated X-vines
$r(\bm{x})=\prod_{e\in E_{1}}r_{a_{e},b_{e}}(x_{a_{e}},x_{b_{e}})\cdot\prod_{j=2}^{q}\prod_{e\in E_{j}}c_{a_{e},b_{e};D_{e}}\left({R_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}}),R_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}})}\right),$ (5.10)
The right-hand side of (5.10) involves $\sum_{j=2}^{q}(d-j)$ bivariate copula densities. The truncation level $q$ allows to tune the trade-off between sparsity and flexibility. If $q=1$, the second product is empty and the model is a Markov tree as in Engelke and Hitz (2020), Segers (2020) and Engelke and Volgushev (2022). Increasing $q$ and adding trees (‘layers’) yields more complex dependence models. Truncated X-vine specifications will be shown at work in Section 9.
@@ 6.1 Inverted multivariate Pareto distributions
The $d$-variate random vector $\bm{Z}$ has an inverted multivariate Pareto distribution if there exists a $d$-variate tail copula measure $R$ such that $\operatorname{\mathsf{P}}(\bm{Z}\in B)=R(B\cap\mathbb{L})/R(\mathbb{L})$, for Borel sets $B$, with $\mathbb{L}=\left\{{\bm{x}\in(0,\infty]^{d}:\min\bm{x}<1}\right\}$; equivalently, if there exists a multivariate Pareto random vector $\bm{Y}$ such that $\bm{Z}=1/\bm{Y}$.
If $R$ is concentrated on $(0,\infty)^{d}$ and has density $r$, then $\bm{Z}$ has probability density $\bm{x}\mapsto r(\bm{x})\operatorname{\mathds{1}}(\bm{x}\in\mathbb{L})/R(\mathbb{L})$. In general, for $j\in[d]$, the conditional distribution of $Z_{j}$ given $Z_{j}<1$ is uniform on $(0,1)$; this follows from the marginal constraint (2.2). For Borel sets $B$, we have
@@
For non-empty $J\subseteq[d]$, we have $\left({t\bm{U}_{J}\mid\min\bm{U}_{J}<1/t}\right)\rightsquigarrow\bm{Z}_{|J}\stackrel{{\scriptstyle d}}{{=}}\left({\bm{Z}_{J}\mid\min\bm{Z}_{J}<1}\right)$, as $t\rightarrow\infty$. The conditional marginal $\bm{Z}_{|J}$ has a $|J|$-variate inverted multivariate Pareto distribution associated with the marginal tail copula measure $R_{J}$ in (2.6). In contrast, $\bm{Z}_{J}$ does not necessarily have an inverted multivariate Pareto distribution, as $\min\bm{Z}_{J}<1$ is not guaranteed.
6.2 Sampling from (inverted) multivariate Pareto distributions
Through equations (2.10) and (6.2), the (inverted) multivariate Pareto distribution serves as a model for a random vector conditionally on the event that at least one variable takes a value far in the tail of its respective marginal distribution. The L-shaped support of the (inverted) multivariate Pareto distribution makes direct random sampling from it a little awkward. Lemma 2 in Engelke and Hitz (2020) provides an ingenious algorithm that reduces the task of sampling a multivariate Pareto random vector $\bm{Y}$ to sampling the conditional distributions $\left({\bm{Y}\mid Y_{j}>1}\right)$ for every $j\in[d]$. Below, we study the equivalent problem of simulating from the conditional distribution of $\left({\bm{Z}\mid Z_{j}<1}\right)$ for X-vine tail copula densities.
We will do so by inverting the Rosenblatt transformation (Rosenblatt, 1952), applying conditional quantile functions with an increasing number of conditioning variables successively to independent uniform random variables. A judicious choice of the ordering of the variables permits to compute the required conditional quantile functions recursively in terms of the bivariate ingredients of the X-vine specification. This sampling order (Cooke et al., 2015) is encoded by the permutation constructed in the next lemma.
@@ 6.2 Sampling from (inverted) multivariate Pareto distributions
Let $\mathcal{V}$ be a regular vine sequence on $d$ elements. For all $j\in[d]$, there exists a permutation $\sigma_{j}$ of $\left\{{1,\ldots,d}\right\}$ such that (i) $\sigma_{j}(1)=j$ and (ii) there exist edges $e_{j,1}\in E_{1},\ldots,e_{j,d-1}\in E_{d-1}$ such that $\sigma_{j}(k)\in\mathcal{C}_{e_{k-1}}$ and $\left\{{\sigma_{j}(i):i=1,\ldots,k}\right\}=A_{e_{j,k-1}}$ for all $k\in\left\{{2,\ldots,d}\right\}$.
Let $r$ be an X-vine tail copula density as in Definition 5.3. For each edge $e\in E_{2}\cup\ldots\cup E_{d-1}$ and for fixed $10<u_{b_{e}}<$ 1, let $u_{a_{e}}\mapsto C_{a_{e}|b_{e};D_{e}}^{-1}\left({u_{a_{e}}|u_{b_{e}}}\right)$ be the inverse of the distribution function $u_{a_{e}}\mapsto C_{a_{e}|b_{e};D_{e}}\left({u_{a_{e}}|u_{b_{e}}}\right)$, with $C_{a_{e}|b_{e};D_{e}}$ defined in (5.3). Similarly for $C_{b_{e}|a_{e};D_{e}}^{-1}$. Recall from Proposition 3.1 that for $i\in[d]$, non-empty $J\subseteq[d]\setminus\left\{{i}\right\}$, and $\bm{x}_{J}\in(0,\infty)^{d}$ such that $r_{J}(\bm{x}_{J})>0$, the quantile function $u_{i}\mapsto R^{-1}_{i|J}(u_{i}|\bm{x}_{J})$ is the inverse of the distribution function $x_{i}\mapsto R_{i|J}(x_{i}|\bm{x}_{J})$. Inverting (5.4) yields the recursive relations
@@
$Z_{j}^{(j)}=W_{j}~{}\text{and}~{}Z_{\sigma_{j}(k)}^{(j)}=R^{-1}_{\sigma_{j}(k)|\sigma_{j}(\left\{{1,\ldots,k-1}\right\})}\left({W_{\sigma_{j}(k)}\mid\bm{Z}_{\sigma_{j}(\left\{{1,\ldots,k-1}\right\})}^{(j)}}\right),\quad k\in\left\{{2,\ldots,d}\right\}.$ (6.4)
@@
The simulation algorithm based on Proposition 6.3 relies on structure matrices encoding regular vine sequences as explained in Appendix A in the supplement. In Section 8, we apply the algorithm to assess the estimation methods from Section 7 through Monte Carlo experiments.
@@ 7 Estimation and model selection
Let $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,d})$ for $i\in\left\{{1,\ldots,n}\right\}$ be an independent random sample from a distribution function $F$ with continuous but unspecified margins $F_{1},\ldots,F_{d}$ and whose survival copula $C$ has lower tail copula $R$ (Definition 2.1). Suppose that the tail copula density $r$ is an X-vine with specification $\left({\mathcal{V},\mathcal{R},\mathcal{C}}\right)$ (Definitions 5.3 and 5.5 and Theorem 5.6). We propose a procedure to estimate $r$ from the excesses over a high multivariate threshold.
@@ 7.1 Copulas and inverted multivariate Pareto distributions
Let $r$ be a $d$-variate tail copula density, not necessarily an X-vine. In Sklar’s theorem (Proposition 3.1), suppose that the copula density $c_{I;J}$ satisfies the simplifying assumption (Definition 3.4). The following proposition shows how to transform an inverted multivariate Pareto random vector $\bm{Z}$ associated to $r$ into a random vector with density $c_{I;J}$.
@@
By statement (ii), the density of $\left({R_{i|J}(Z_{i}|\bm{Z}_{J})}\right)_{i\in I}$ given $\bm{Z}_{J}\in A$ is equal to $c_{I;J}$ for any non-empty set $A\subseteq\left\{{\bm{z}\in(0,\infty)^{J}:\min\bm{z}<1}\right\}$. For estimation, we will use this property for $A=(0,1)^{J}$, requiring in effect that all variables (rather than at least one) in $J$ exceed a high threshold. We do so in order to avoid a potential bias stemming from including too many non-extreme values in the procedure. An alternative would be to opt for a censored likelihood approach (Ledford and Tawn, 1996).
@@ 7.2 Sequential maximum likelihood estimation of X-vines
Let the $d$-variate tail copula density $r$ be an X-vine specified by $(\mathcal{V},\mathcal{R},\mathcal{C})$ as in Theorem 5.6. Assume that the bivariate tail copula densities ($\mathcal{R}$) and bivariate copula densities ($\mathcal{C}$) belong to specified parametric families. Let the parameter vector be denoted by $\bm{\theta}=(\theta_{\mathcal{R}},\theta_{\mathcal{C}})$: here, $\theta_{\mathcal{R}}=(\theta_{e})_{e\in E_{1}}$ contains the parameters (or parameter vectors) $\theta_{e}\in\Theta_{e}$ associated with each pairwise tail copula density $r_{a_{e},b_{e}}$ for $e\in E_{1}$, while $\theta_{\mathcal{C}}=\left({\theta_{\mathcal{C},j}}\right)_{j=2}^{d-1}$ for $\theta_{\mathcal{C},j}=(\theta_{e})_{e\in E_{j}}$ denotes the parameters (or parameter vectors) $\theta_{e}\in\Theta_{e}$ associated with each bivariate copula density $c_{a_{e},b_{e};D_{e}}$ for $e\in\bigcup_{j=2}^{d-1}E_{j}$. While it is possible to derive the full likelihood of an X-vine (inverted) multivariate Pareto distribution, performing parameter estimation with the full model in high dimensions is challenging. Instead, using the X-vine decomposition into bivariate components and recursively defined quantities (Theorem 5.6), we outline a sequential procedure for parameter estimation, tree by tree. This approach is inspired by the one for regular vine copulas (see, e.g., Czado, 2019), but with suitable adaptations to the extreme value context.
(1) Standardising the margins and selecting sub-samples.
Recall that $\bm{X}_{1},\ldots,\bm{X}_{n}$ is an independent random sample from $F$, with survival copula $C$, tail copula $R$ and tail copula density $r$. For $i=1,\ldots,n$ and $j\in[d]$, let $\widehat{U}_{i,j}=1-\widehat{F}_{j}(X_{i,j})$, where $\widehat{F}_{j}$ denotes any estimator of the marginal distribution function $F_{j}(x)=\operatorname{\mathsf{P}}(X_{i,j}\leqslant x)$. One possibility is the empirical distribution function, and to avoid boundary effects, we set
@@
By Eq. (6.2), for large $t>0$, the rescaled points $t\widehat{\bm{U}}_{i}$ for $i\in\left\{{1,\ldots,n}\right\}$ such that $\min\widehat{\bm{U}}_{i}<1/t$ constitute pseudo-observations from a distribution that approximates the inverted multivariate Pareto distribution associated with $r$. We set $t=n/k$ where, in an asymptotic setting, $k=k_{n}\in\left\{{1,\ldots,n}\right\}$ satisfies $k\to\infty$ and $k/n\to 0$. For $j\in[d]$ and $\widehat{U}_{i,j}$ as in (7.1), let
@@
be the set of indices $i$ corresponding to the $k$ largest observations for the $j$th component. Further, put $K_{J}=\bigcap_{j\in J}K_{j}$ and $K=\bigcup_{j\in[d]}K_{j},$ which are, respectively, the set of indices $i$ corresponding to large observations in all variables in $J\subseteq[d]$ simultaneously and the set of indices with large observations in at least one variable. Write $\widehat{\bm{Z}}_{i}=(\widehat{Z}_{i,j})_{j\in[d]}$ where $\widehat{Z}_{i,j}=\left({n/k}\right)\widehat{U}_{i,j}$. In view of Eq. (6.2), we treat $\{\widehat{\bm{Z}}_{i}\}_{i\in K}$ as a sample of $|K|$ pseudo-observations of the inverted multivariate Pareto distribution associated with $r$. The sample size $|K|$ is random, and from (7.1), we have $k\leqslant|K|\leqslant dk$. For non-empty $J\subseteq[d]$ and $i\in K$, we write $\widehat{\bm{Z}}_{i,J}=(\widehat{Z}_{i,j})_{j\in J}$.
(2) Estimating the tail copula parameters $\theta_{\mathcal{R}}$ in $T_{1}$.
For each edge $e=\left\{{a_{e},b_{e}}\right\}\in E_{1}$, we estimate the parameter (vector) $\theta_{e}$ associated with the bivariate tail copula density $r_{a_{e},b_{e}}(\,\cdot\,;\theta_{e})$. By Eq. (6.1), for $j\in\left\{{a_{e},b_{e}}\right\}$, the conditional density of $(Z_{a_{e}},Z_{b_{e}})$ given $Z_{j}<1$ is $r_{a_{e},b_{e}}(z_{a_{e}},z_{b_{e}};\theta_{e})\operatorname{\mathds{1}}(z_{j}<1)$. We use maximum pseudo-likelihood estimation to fit this density to the sub-samples $(\widehat{Z}_{i,a_{e}},\widehat{Z}_{i,b_{e}})$ for both $i\in K_{a_{e}}$ and $i\in K_{b_{e}}$, with $K_{j}$ as in Eq. (7.2). More precisely, we maximise each of the two pseudo-likelihoods
@@
over $\theta_{e}^{(j)}\in\Theta_{e}$, yielding estimates $\hat{\theta}_{e}^{(a_{e})}$ and $\hat{\theta}_{e}^{(b_{e})}$, respectively. The final estimate is $\hat{\theta}_{a_{e},b_{e}}=\{\hat{\theta}_{e}^{(a_{e})}+\hat{\theta}_{e}^{(b_{e})}\}/2$. The idea of averaging maximum pseudo-likelihood estimators on product spaces has already been proposed for the Hüsler–Reiss model (Engelke et al., 2015; Engelke et al., 2022). Other estimation approaches include censored likelihoods (Ledford and Tawn, 1996; de Haan et al., 2008) or empirical stable tail dependence functions (Einmahl et al., 2008, 2018).
(3) Estimating the copula parameters $\theta_{\mathcal{C}}$ in $T_{2},\ldots,T_{d-1}$.
Estimation of the parameters associated with edges $e=(a_{e},b_{e};D_{e})$ in $E_{j}$ for $j\in\left\{{2,\ldots,d-1}\right\}$ is based on a similar procedure as the one employed in regular vine copulas (Czado, 2019, Section 7.2). The main difference concerns the definition of the pseudo-observations: if $\hat{\theta}(E_{1:(j-1)})$ denotes the parameter estimates associated with the edges in $E_{1:(j-1)}=E_{1}\cup\cdots\cup E_{j-1}$, by Proposition 7.1(i), for $i\in K_{D_{e}}$, pseudo-observations $(\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}})$ from $c_{a_{e},b_{e};D_{e}}$ can be defined by
$\widehat{U}_{i,a_{e};D_{e}}=R_{a_{e}|D_{e}}\left({\widehat{Z}_{i,a_{e}}\mid\widehat{\bm{Z}}_{i,D_{e}};\;\hat{\theta}(E_{1:j-1})}\right),\quad\widehat{U}_{i,b_{e};D_{e}}=R_{b_{e}|D_{e}}\left({\widehat{Z}_{i,b_{e}}\mid\widehat{\bm{Z}}_{i,D_{e}};\,\hat{\theta}(E_{1:j-1})}\right).$ (7.4)
The full procedure is detailed in Section D of the supplement.